Find the equilibrium price and commodity.
Pd= 200-5Q
Ps= 92+4Q
Can Price be expressed as this function? If yes, please explain with graph and theory.
1
Expert's answer
2018-08-24T07:58:08-0400
Pd= 200-5Q, Ps= 92+4Q.
In equilibrium Qd = Qs and Pd = Ps, so:
200 - 5Q = 92 + 4Q,
9Q = 108,
Q = 12 units,
P = 200 - 5*12 = 140..
Price can be expressed as this function.